### v1

# Monitoring-feedback induced entanglement relaxations in a tilted free fermionic chain

Abstract
Recent years have witnessed measurement-induced entanglement phase transitions attracting significant attention, yet the feedback-induced skin effect in tilted fields remains understudied. This work investigates entanglement relaxation and dynamical transitions in a fermionic chain subject to generalized monitoring and an applied tilted field. We show that the feedback-induced skin effect persists with tilting, though relaxation time increases. For small tilts, bipartite entanglement exhibits a widened metastable phase, transitioning from a logarithmic-law to an area-law phase. In contrast, strong tilts accelerate relaxation, leading to a single area-law phase. Moreover, due to the unique properties of the tilted chain, we examine the impact of the boundary on entanglement relaxation. This work not only deepens our understanding of the role of feedback in measurement-induced entanglement phase transitions but also enhances our comprehension of the non-Hermitian skin effect beyond the no-click limit.
preprint: APS/123-QED
I Introduction
Depicting the evolution of quantum matter far from equilibrium with universal quantum dynamics might be notoriously challenging. However, considerable efforts to understand out-of-equilibrium dynamics in lattice models [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11], randum unitary circuits [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], black hole information theory [23, 24, 25], and quantum field theory [26, 27, 28] reveal that universal patterns may emerge in the entanglement structures of many-body systems. The core issues in condensed matter physics today, such as quantum thermalization [29, 30, 31], quantum chaos [32, 33, 34], holography [28, 35], and information scrambling [36, 37, 38], all live in the shadow of these entanglement patterns evolution.
In unitary entanglement relaxation dynamics (hereafter refer to entanglement relaxation), local perturbations spread entanglement within the Lieb-Robinson light cone [39], initially increasing entanglement entropy until it saturates. Growth patterns vary by system type: integrable systems show slower growth [40, 41, 42, 43, 44], non-integrable systems exhibit faster thermalized growth [45, 46], and many-body localized systems display logarithmic growth due to localization effects [47, 46, 48]. Interestingly, introducing weak measurements leads to a competition between unitary and non-unitary dynamics in the system, resulting in diverse entanglement relaxation patterns as well, thereby driving measurement-induced entanglement phase transitions (MIPT) [12, 13, 14, 49, 50, 51, 52, 53, 18].
On the one hand, in the no-click limit [54], MIPT corresponds to purely non-Hermitian dynamics, where the post-selection mechanism of measurement-feedback effectively erases trajectories with quantum jumps. This phenomenon often directly reduces to an entanglement dynamics problem involving a Hatano-Nelson type [55, 56, 57] effective Hamiltonian, where entanglement relaxation is accompanied by a steady-state entanglement phase transition, eventually equilibrating to a non-Hermitian skin state [58, 59]. The entanglement relaxation and MIPT without quantum jumps have been extensively studied in free lattice models[60, 61, 11, 62, 63, 64], even including its interaction with localization potentials [60, 61, 65, 11]. However, implementing post-selection experimentally is notably difficult. On the other hand, entanglement relaxation and MIPT under quantum jumps [66] can also be described through stochastic dynamics in quantum many-body trajectories. In this context, if the effective Hamiltonian exhibits many-body skin pumping, quantum jump noise drives the entanglement steady state into a highly entangled configuration, rather than allowing it to settle into a low-entanglement non-Hermitian skin state. Consequently, the information scrambling from quantum jumps makes it challenging to observe the many-body skin effect in experimental settings.
However, substantial recent efforts, have been made to understand entanglement relaxation with quantum jumps under generalized monitoring [67], especially within the context of the skin effect, in lattice models and quantum circuits [68, 69, 67, 70, 71]. There, feedback operations suppress the dynamical noise caused by quantum jumps, allowing the system’s long-time steady-state to be dominated by the skin effect of the non-Hermitian effective Hamiltonian. Under the rescaled time ${\rm time}/L$, where $L$ is the system size, the feedback-induced skin effect drives non-monotonic bipartite entanglement growth. Even more surprisingly, this process reveals a novel dynamical transition that exhibits robustness against disordered-induced localization perturbations [70]. Nevertheless, whether this emergent entanglement dynamical transition induced by monitoring-feedback exists in disorder-free localized systems, particularly in systems with a Wannier-Stark (or tilted) potential [72], has not been studied before. The Wannier-Stark potential not only leads to hyperexponentially localized wave functions distinct from disorder in single-particle systems but also offers significant numerical and theoretical advantages, as it eliminates the need for disorder averaging. Additionally, with the rise of quantum simulation experimental platforms [73, 74], discussions on the non-equilibrium behavior of quantum many-body systems are no longer confined to Hilbert space, and disorder-free localization systems offer promising experimental prospects. A natural question that arises is: what is the fate of feedback-induced skin effects in the entanglement relaxation under the Wannier-Stark potential?
In this work, we analyze monitoring-feedback induced entanglement relaxation in a tilted free fermionic chain using the quantum trajectory method. Averaging over multiple Stochastic Schrödinger Equation (SSE) evolutions with quantum jumps on the time scale ${\rm time}/L$, we observe bipartite entanglement dynamics under a Wannier-Stark potential. Initially, entanglement enters a transient metastable phase, followed by long-time relaxation to a low-entanglement steady state, akin to the skin effect. With minor tilting, non-monotonic entanglement relaxation and dynamical transition persist. However, strong tilting appears to accelerate relaxation, potentially suppressing the entanglement transition. Additionally, we examine boundary conditions, highlighting insights into feedback-monitored entanglement dynamics and the non-Hermitian skin effect with quantum jumps. Our findings not only shed light on the entanglement relaxation dynamics of feedback-monitored quantum many-body systems under the competition of disorder-free localization but also contribute to a deeper understanding of the non-Hermitian skin effect with quantum jumps.
This paper is organized as follows. Section II introduces the model along with the monitoring-feedback protocol. In Section III, we present the analysis of trajectory-averaged relaxations, focusing on the time evolution of the average density distribution. Section IV provides numerical evidence of the entanglement dynamical transitions under varying tilts. Additionally, we explore the quasiparticle interpretation of the bulk and edge dynamics in Section V. Finally, Section VI offers a summary and outlines future research directions. Additional numerical calculation data and discussions are included in the Appendix.
II Setups
In this work, we consider a tilted free fermion chain of length $L$ and choose open boundary conditions (OBCs) unless otherwise specified. The system is described by the following Hamiltonian,
$\hat{H}=\sum_{l}^{L-1}{J(\hat{c}_{l}^{\dagger}\hat{c}_{l+1}+H.c.)}+\Delta\sum_{l}^{L}{l\hat{c}_{l}^{\dagger}\hat{c}_{l}},$ (1)
where $\hat{c}_{l}^{\dagger}$ ($\hat{c}_{l}$) is the fermion creation (annihilation) operator at site $l$, $J$ is the strength of nearest-neighbor hopping which is set to $1$ hereafter for convenience, and $\Delta$ is the strength of the tilted potential, i.e., the on-site potential increases linearly with the site index $l$ with the potential difference between the nearest-neighbor sites being $\Delta$. Introducing the continuous measurements, the evolution of the open quantum system is described by the Lindblad master equation which is a formidable challenge for numerical simulation. Thus, to simply this process, we instead consider the SSE [66, 75, 76, 77, 78, 79] corresponding to the trajectory evolution of a pure state $\left|\psi(t)\right\rangle$
$d\left|\psi(t)\right\rangle$ $\ =-i\hat{H}_{\text{eff}}\left|\psi(t)\right\rangle dt$ (2)
$\ +\sum_{l}^{L-1}\left(\frac{\hat{L}_{l}\left|\psi(t)\right\rangle}{\sqrt{\left\langle\hat{L}_{l}^{\dagger}\hat{L}_{l}\right\rangle}}-\left|\psi(t)\right\rangle\right)dW_{l},$
where $\hat{L}_{l}$ is a quantum jump operator induced by weak measurement, $\langle\hat{O}\rangle=\left\langle\psi(t)\right|\hat{O}\left|\psi(t)\right\rangle$ is the expectation value of operator $\hat{O}$, $dW_{l}$ is an independent and discrete Poisson random variable, which takes values from {0, 1} with mean value $\langle\hat{L}_{l}^{\dagger}\hat{L}_{l}\rangle\gamma dt$. $\hat{H}_{\text{eff}}$ is non-Hermitian effective Hamiltonian, $\hat{H}_{\text{eff}}=\hat{H}-i\gamma/{2}{\textstyle\sum_{l}^{L-1}{\hat{L}_{l}^{\dagger}\hat{L}_{l}}}$, where $\gamma$ is the monitoring rate.
Figure 1: (a) The influence of the continuous measurement with the feedback under the tilted potential. (b) OBCs, the dynamical phase diagram of the trajectory averaged bipartite entanglement entropy in the plane (${\rm time}/L$, $\Delta$) with $L=192$. The figure is primarily segmented by log-law region, area-law region, and no exact edge.
We consider the continuous measurements with feedback, inspired by Refs. [67]. The quantum jump operator is written as
$\hat{L}_{l}=e^{i\theta\hat{n}_{l+1}}\hat{d}_{l}^{\dagger}\hat{d}_{l},$ (3)
where $e^{i\theta\hat{n}_{l+1}}$ is the feedback (see Appendix B for the feedback protocol considering the first site under PBCs), $\hat{n}_{l+1}$ is the particle number operator, and $\hat{d}^{\dagger}_{l}=(\hat{c}^{\dagger}_{l}-i\hat{c}^{\dagger}_{l+1})/\sqrt{2}$ is the creation operator of a right-moving quasimode. If $\theta\neq 0$, the feedback will convert the right-moving quasimode to the left-moving one, and we set $\theta=\pi$ without loss of generality. Based on the previous text, we now can write $\hat{H}_{\text{eff}}$ [see FIG. 1 (a)] as
$\hat{H}_{\text{eff}}$ $\ =\sum_{l}^{L-1}{[(1+\frac{\gamma}{4})\hat{c}_{l}^{\dagger}\hat{c}_{l+1}+(1-\frac{\gamma}{4})\hat{c}_{l+1}^{\dagger}\hat{c}_{l}]}$ (4)
$\ -i\frac{\gamma}{4}\sum_{l}^{L-1}{(\hat{n}_{l}+\hat{n}_{l+1})}+\Delta\sum_{l}^{L}{l\hat{n}_{l}}.$
In the subsequent sections, we set $\gamma=0.5$, which corresponds to the case of the critical entanglement phase caused by the weak measurement.
Figure 2: The trajectory averaged density distribution evolution with respect to rescaled time for $L=128$. (a) OBC, $\Delta=0$; (b) OBC, $\Delta=0.6$; (c) PBC, $\Delta=0$; (d) PBC, $\Delta=0.6$.
We choose the Neél state as the initial state, i.e., $\left|\psi(t=0)\right\rangle=\left|0101...01\right\rangle$. Since both $\hat{H}_{\text{eff}}$ and $\hat{L}_{l}$ are quadratic, we can utilize the Gaussian state to efficiently simulate each trajectory evolution of $\left|\psi(t)\right\rangle$. We note the observables are averaged 200-800 trajectories, represented by the variable with an overline (e.g., trajectory averaged density distribution $\overline{\langle\hat{n}_{l}\rangle}$). Moreover, the observables can be easily obtained from the correlation matrix $C_{ij}=\langle\hat{c}^{\dagger}_{i}\hat{c}_{j}\rangle$. $\langle\hat{n}_{l}\rangle$ are the diagonal elements of $C_{ij}$. For the trajectory averaged bipartite entanglement entropy $\overline{S}_{L/2}$  [80, 81, 82, 83, 84], we first choose the left half chain as the subsystem, then get sub correlation matrix $C_{ij}^{L/2}$ from $C_{ij}$ and lastly obtain ${S}_{L/2}$ based on all eigenvalues $\{\lambda_{i}\}$ of $C_{ij}^{L/2}$ with $S_{L/2}=-{\textstyle\sum_{i}{\lambda_{i}\log{(\lambda_{i})}}+(1-\lambda_{i})\log{(1-\lambda_{i})}}$. The trajectory-averaged velocity of particles, $\overline{v}$, as derived from the effective Hamiltonian $\hat{H}_{\text{eff}}$ and the correlation function $C_{ij}$, reflects the dynamics of entanglement relaxation [70]. See Appendix A for more details.
III Trajectory averaged relaxations
Figure 3: (a) and (c) show the time evolution of the trajectory averaged bipartite entanglement entropy $\overline{S}_{L/2}$ for different $L$ under OBCs and PBCs separately. The green solid lines represent the absence of the tilted potential ($\Delta=0$), meanwhile, the orange solid lines represent the presence of the tilted potential ($\Delta=0.6$). The colors ranging from light to dark correspond to $L=64,96,128,160,192$. In the zoomed inset, the dashed black highlights the transition point, which separates the log-law and area-law regions. Inset: $\overline{S}_{L/2}$ versus $\log{L}$ at ${\rm time}/L=$ $0.5$ (diamond), $1.8$ (square). (b) and (d) illustrate the maximum bipartite entanglement entropy $\overline{S}_{L/2}^{max}$ with respect to $\Delta$ for $L=64,96,128,160,192$ (from light to dark colors). Inset: $\overline{S}_{L/2}^{max}$ versus $\log{L}$ for different $\Delta$.
We first investigate how the time-evolving average density distribution $\overline{\langle\hat{n}_{l}\rangle}$, behaves under different boundary conditions: open boundary conditions (OBCs) and periodic boundary conditions (PBCs). Additionally, we will examine the effects of the presence and absence of a tilted potential. In FIG. 2 (a), (b), and (d), the late-time state is characterized by a notable feature: particles predominantly occupy the left half of the chain, approaching what is commonly referred to as a domain wall state however without a distinct boundary. The phenomenon observed is known as the feedback-induced skin effect, as detailed in Ref. [67]. Specifically, under OBCs, a system under the tilted potential with a strength of $\Delta=0.6$ exhibits a prolonged residence in the non-equilibrium state compared to a system without the tilted potential ($\Delta=0$). This extended duration in the non-equilibrium state is likely attributable to the restriction of the tilted potential for the particle motion.
FIG. 2 (c) stands out from the others in the series, as it demonstrates that $\overline{\langle\hat{n}_{l}\rangle}$ remains relatively uniform across all lattice sites for the majority of the observation period. This uniformity can be attributed to the absence of edges and the dynamics are determined by the bulk thus preventing the skin effect. Furthermore, when comparing FIG.  2 (c) and (d), despite both under PBCs, the system in (d) does not exhibit such uniform distribution. Conversely, as time evolves the relaxation process closely mirrors the behavior observed in (b) and late-time $\overline{\langle\hat{n}_{l}\rangle}$ is mainly distributed on the left of the chain. The feedback-induced skin effect still exists in this system despite under PBCs. The reason is that the presence of the tilted potential creates a significant disparity between the first and the last on-site potentials, which generates a pseudo edge. Therefore, the particles hardly move from the first to the last site. Under PBCs, the pseudo edge plays a role similar to that of the edge under OBCs, restricting the bulk dynamics of particles. The dashed line shown in FIG. 2 indicates a transition point within the relaxation process, that will be further discussed in the following section.
IV Entanglement relaxations and dynamical transition
Now, we delve into the dynamics of entanglement relaxation by examining the bipartite entanglement entropy, denoted as $\overline{S}_{L/2}$, across various $L$ and $\Delta$. Under OBCs, in FIG. 3 (a), $\overline{S}_{L/2}$ first sharply increases, then after the transient metastable value, slowly declines until the steady value irrespective of the presence or absence of the tilted potential. The entanglement relaxation with the tilted potential ($\Delta=0.6$) requires more time than the case without it ($\Delta=0$). Notably, there is a cross point during the decline of $\overline{S}_{L/2}$ for different $L$ and the same $\Delta$. We emphasize them with dashed lines in zoomed insets, called transition points. The transition point separates the log-law and area-law regions. From the example of FIG. 3 (a) inset, in the left of the transition point the relationship between $\overline{S}_{L/2}$ and $\log{L}$ is a linear increase; but in the right, $\overline{S}_{L/2}$ remains constant regardless of $\log{L}$. For $\Delta=0$, separated time (rescaled time ${\rm time}/L$) is about $0.79$, which is consistent with the results in Ref. [70]; for $\Delta=0.6$, the time is approximately $1.47$. As the duration of the entanglement-relaxation process increases, the transition point is delayed, corresponding to the zone below the no exact edge shown in FIG. 1 (b) which are separated from the area law region with a fuzzy boundary. We have noticed that if $\Delta$ is significant enough and entanglement relaxations are prolonged enough, there might be no dynamical entanglement phase transition. For this case, $\overline{S}_{L/2}$ stays in the area-law region after an initial rapid increase. This novel phenomenon is demonstrated in the zone located above the non-exact edge in FIG. 1 (b).
Figure 4: Under OBCs, the dynamical phase diagram of $\overline{S}_{L/2}$ in the plane $\left(\text{time}/L,\Delta\right)$ for $L=192$ for relatively weak $\Delta$.
In FIG. 3 (c), considering $\Delta=0$ under PBCs, after a brief increase, $\overline{S}_{L/2}$ remains stable, which is log-law for the steady value. From the inset, $\overline{S}_{L/2}$ increases linearly with respect to $\log{L}$, and different-time relationships are basically the same. If $\Delta=0.6$, the behavior of $\overline{S}_{L/2}$, starts with a sharp increase, moves to metastable value, and finally declines to a steady value, which is the same as the case under OBCs, in terms of the entanglement relaxations and entanglement transitions due to the pseudo edge.
The maximum trajectory averaged bipartite entanglement entropy $\overline{S}_{L/2}^{max}$ could be found in metastable value. FIG. 3 (b) and (d) illustrate how $\overline{S}_{L/2}^{max}$ changes with $\Delta$ for different $L$, under OBCs and PBCs separetely. In FIG. 3 (b), $\overline{S}_{L/2}^{max}$ increases with $\Delta$ until the peak, about $0.4$, then decreases. $\overline{S}_{L/2}^{max}$ undergoes a transition from log law to area law, and we provide some illustrative instances of the different linear relationships between the $\overline{S}_{L/2}^{max}$ and $\log{L}$ in the inset of FIG. 3 (b). FIG. 3 (d) shows the PBCs case, we find that only $\Delta=0$ there is a huge difference between PBCs and OBCs because of the absence of the pseudo edge. We can also see the transition from log law to area law as $\Delta$ increases.
We have constructed the dynamical phase diagram depicted in FIG. 4 by studying $\overline{S}_{L/2}$ at various times and across different values of $\Delta$. This figure can be viewed as an enlarged detail of FIG. 1 (b), specifically focusing on the low $\Delta$ zone. In FIG. 3 (a), the entanglement relaxation processes described by the darkest green and orange solid lines can also be found in FIG. 4. Upon observation, it is evident that the duration of entanglement relaxation extends as $\Delta$ increases. Concurrently, the transition point that distinguishes between the log-law and area-law regions remains consistent but shifts to the right. The postponed transition point can be attributed to the fact that the tilted potential decelerates the motion of the particles. When the motion is sufficiently slowed, the outcome is depicted in FIG. 1(b).
Figure 5: $L=64$ and OBC. (a) shows the variation of the trajactory averaged $\overline{v}$ over rescaled time. (b) illustrates that the relaxtion acceration $\overline{RA}$ versus $\Delta$. The data of $\overline{RA}$ are from the linear fit of the decaying $\overline{v}$ in (a).
V The interplay between bulk dynamics and edge effect
From the perspective of particle dynamics, we conduct discussions of entanglement relaxation. Under OBCs, we have chosen the trajectory averaged velocity of particles, represented as $\overline{v}$ [70], as our key observable. In FIG. 5 (a), particles are initially given a velocity directed to the left. This velocity then decays linearly over time, eventually to stable zero, which marks the end of the relaxation process. Subsequently, the particles will no longer move. The decay of $\overline{v}$ is primarily due to the edge effect. Furthermore, as $\Delta$ increases, the initial velocity of the particles decreases, resulting in a longer relaxation time required for the system to return to equilibrium. This indicates that $\Delta$ holds the bulk dynamics of the particles.
We have determined the slope of the decay for $\overline{v}$ for different $L$ and depicted these values in FIG. 5 (b), which represent the averaged relaxation acceleration, denoted as $\overline{RA}$. It is observed that $\overline{RA}$ decreases as $\Delta$ increases. The phenomenon can be explained by the fact that the smaller the initial velocity of the particles, the longer time it takes for them to reach the edge, and the edge effects exert less influence. In other words, the tilted potential governs the bulk dynamics and influences the relaxation time to equilibrium.
VI Conclusion and outlook
In summary, we examine the entanglement relaxation dynamics induced by generalized monitoring in a tilted free fermion chain under OBCs and PBCs. Under small tilts, for trajectory averaged bipartite entanglement entropy $\overline{S}_{L/2}$ in rescale time ${\rm time}/L$, there is an entanglement phase transition separating the log-law (induced by weak measurement) and area-law (induced by feedback-driving skin effect) regions and such behavior will be delayed as the increase of $\Delta$. However, under strong tilts, only the area-law region exists, and there may be no exact transition behavior, regardless of how long ${\rm time}/L$ is. On the other hand, the tilted potential can result in the pseudo edge for PBCs, which makes the trajectory averaged density distribution $\overline{\langle\hat{n}_{l}\rangle}$ predominantly concentrate on the left half of the chain just like corresponding condition under OBCs which should have been uniformly distributed across all sites. As long as there is a certain degree of the tilted potential, $\overline{S}_{L/2}$ performs the same under OBCs and PBCs. Thus, the pseudo edge under PBCs and the edge under OBCs play the same role in the entanglement relaxation. Edge effects influence the motion of particles, causing the velocity of particles decreases uniformly. The tilted potential restricts the initial velocity of particles. The interplay between them can be concluded that the smaller the initial velocity of the particles, the later most particles reach the edge, and the weaker the influence of boundary effects. Moreover, the quantum jump with the edge feedback could restrict the pseudo-edge effect to some extent under PBCs.
These findings not only expand the understanding of how feedback influences measurement-induced entanglement phase transitions but also enhance the knowledge of the non-Hermitian skin effect [85] with quantum jumps. In the future, it will be interesting to both re-examine various types of non-Hermitian skin effects driven by feedback monitoring, such as the critical non-Hermitian skin effect [86, 87] without quantum jumps, higher-order skin effects [88], and occupation-particle-dependent skin effects [89], and to explore a universal understanding of feedback-monitored entanglement relaxation.
Acknowledgements.
We thank the helpful discussions with Ze-Chuan Liu and Yu-Peng Wang. JXZ acknowledges the National Natural Science Foundation of China (Grant No. 11874316 and No.12374046), the National Basic Research Program of China (Grant No. 2015CB921103), and the Program for Changjiang Scholars and Innovative Research Team in University (Grant No. IRT13093).
Appendix A Numerical Method
A.1 Numerical simulation for evolution of the state
We currently employ Eq.(2) to evolve $\left|\psi(t)\right\rangle$ treated as a Slater determinant state. We can depict $\left|\psi(t)\right\rangle$ with a matrix $U\left(t\right)$ of size $L\times N$ at time $t$,
$\left|\psi(t)\right\rangle=\prod_{n=1}^{N}{\left(\sum_{l=1}^{L}{[U(t)]_{ln}}\hat{c}_{l}^{\dagger}\right)}\left|0\right\rangle,$ (5)
where $L$ is the system size and $N$ is the number of particles, which satisfy $N=L/2$. $\left|0\right\rangle$ stands for vacuum state. Moreover, any column $U_{n}$ of $U\left(t\right)$ can represent the $n$-th single-particle state,
$\left|\psi_{n}\right\rangle=\sum_{l=1}^{L}{\left([U_{n}]_{l}\hat{c}_{l}^{\dagger}\right)\left|0\right\rangle}.$ (6)
After getting the matrix form $H_{\text{eff}}$ of effective Hamiltonian $\hat{H}_{\text{eff}}$, non-Hermitian evolution can be written as
$U\left(t+\delta t\right)=e^{-iH_{\text{eff}}\delta t}U\left(t\right).$ (7)
Following the evolution, we require the application of QR decomposition to ensure the orthogonality among different single-particle states and the normalization of each individual single-particle state.
In this section, we use the site index $m$ to replace the index $l$ which does not take into consideration the $L$-th site. Now we consider the continuous measurement with feedback and numerical simulation the quantum jump, i.e., $\left|\psi\left(t\right)\right\rangle\longrightarrow{\hat{L}_{m}\left|\psi\left(t\right)\right\rangle}/{\sqrt{\langle\hat{L}_{m}^{\dagger}\hat{L}_{m}\rangle}}$. The quantum jump for probability can be written as $p_{m}=\langle\hat{L}_{m}^{\dagger}\hat{L}_{m}\rangle\gamma\delta t$. Using Wick theorem, i.e., $\langle\hat{d}_{m}^{\dagger}\hat{d}_{m}\hat{d}_{m}^{\dagger}\hat{d}_{m}\rangle=\langle\hat{d}_{m}^{\dagger}\hat{d}_{m}\rangle\langle\hat{d}_{m}^{\dagger}\hat{d}_{m}\rangle+\langle\hat{d}_{m}^{\dagger}\hat{d}_{m}\rangle\langle\hat{d}_{m}\hat{d}_{m}^{\dagger}\rangle$, and fermionic anticommutation relations, i.e., $\{\hat{c}_{i},\hat{c}_{j}^{\dagger}\}=\delta_{ij}$, we get the probability of quantum jump, $p_{m}=\langle\hat{d}_{m}^{\dagger}\hat{d}_{m}\rangle\gamma\delta t=\left\langle\psi(t)\right|\hat{d}_{m}^{\dagger}\hat{d}_{m}\left|\psi(t)\right\rangle\gamma\delta t=|\hat{d}_{m}\left|\psi(t)\right\rangle|^{2}\gamma\delta t$. Note that the linear expansion of $\hat{d}_{m}$ can be written as $\hat{d}_{m}={\textstyle\sum_{l}^{L}{[d_{m}]_{l}\hat{c}_{l}}}$, where $d_{m}$ is a row vector. To simulate the quantum jump, we can find a single-particle state which satisfies $\hat{d}_{m}\left|\psi_{n}\right\rangle\neq 0$ and $n>1$, then move the corresponding column $U_{n}$ to the first due to gauge freedom and transform the rest columns of $U(t)$ by
$U_{n}^{\prime}=U_{n}-\frac{d_{m}U_{n}}{d_{m}U_{1}}U_{1}.$ (8)
The result is
$U^{\prime}\left(t\right)=\begin{bmatrix}d_{m}^{\dagger}&U_{2}^{\prime}&U_{3}^{\prime}&\cdots&U_{N}^{\prime}\end{bmatrix}.$ (9)
Finally, we use the QR decomposition for $e^{i\pi\mathcal{N}_{m+1}}U^{\prime}(t)$, and $\mathcal{N}_{m+1}$ is the matrix form of particle number operator. It should be emphasized that, in our numerical simulation, $\delta t=0.01$ , $l=1,2,3,...,L$, $m=1,2,3,...,L-1$, $n=1,2,3,...,N$, except where other noted. Only with a sufficiently small time step can we better simulate the impact of continuous measurements on the evolution of the state.
A.2 Trajectory averaged observables
The gaussian state allows for the extraction of observables from the correlation matrix $C_{ij}=\langle\hat{c}_{i}^{\dagger}\hat{c}_{j}\rangle=\left[U(t)U^{\dagger}(t)\right]^{T}_{ij}$. Therefore, density distribution $\left\langle\hat{n}_{l}\right\rangle$ can be obtained through the diagonal elements of the correlation matrix.
When calculating the bipartite entanglement entropy $S_{L/2}$, we choose the sub system $[1,L/2]$ and then get the all eigenvalues $\lambda_{i}$} of sub-correlation matrix $C_{ij}^{L/2}$. $S_{L/2}$ is given by  [80, 81, 82, 83, 84]
$S_{L/2}=-\sum_{i=1}^{L/2}{\lambda_{i}\log{(\lambda_{i})}+(1-\lambda_{i})\log{(1-\lambda_{i})}}.$ (10)
Figure 6: PBCs, the trajectory averaged density distribution evolution with respect to rescaled time for $L=128$ and the time step is $0.02$ (a) previous feedback, $\Delta=0$; (b) edge feedback, $\Delta=0$; (c) previous feedback, $\Delta=0.6$; (d) edge feedback, $\Delta=0.6$.
When calculating the velocity of particles under OBCs, taking into account that the nearest-neighbor distance is normalized, we can get the average position for particles depicted by $\langle\hat{x}\rangle={\textstyle\sum_{l}^{L}l\langle\hat{n}_{l}\rangle}/N$. Furthermore, the average velocity per particle $v$ equals $d\langle\hat{x}\rangle/dt$ The master equation,
$\frac{d\rho_{t}}{dt}=-i\hat{H}_{\text{eff}}\rho_{t}+i\rho_{t}\hat{H}_{\text{eff}}^{\dagger}+\gamma\sum_{m}^{L-1}{\hat{L}_{m}\rho_{t}\hat{L}_{m}^{\dagger}},$ (11)
which governs the evolution of the density matrix $\rho_{t}$. Using $\langle\hat{n}_{l}\rangle$ to replace $\rho_{t}$, we have
$\frac{d\left\langle\hat{n}_{l}\right\rangle}{dt}=-i\left\langle\hat{n}_{l}\hat{H}_{\text{eff}}\right\rangle+i\left\langle\hat{H}_{\text{eff}}^{\dagger}\hat{n}_{l}\right\rangle+\gamma\sum_{m}^{L-1}{\left\langle\hat{L}_{m}^{\dagger}\hat{n}_{l}\hat{L}_{m}\right\rangle}.$ (12)
Figure 7: PBCs, the trajectory averaged bipartite entanglement entropy $\overline{S}_{L/2}$ evolves with respect to rescaled time, for different feedback and $\Delta$.
Finally, we employ Wick theorem and get the average velocity per particle,
$v$ $\ =\frac{1}{N}\sum_{l}{i\left(l\left\langle\hat{H}_{\text{eff}}^{\dagger}\hat{n}_{l}\right\rangle-l\left\langle\hat{n}_{l}\hat{H}_{\text{eff}}\right\rangle\right)}$ (13)
$\ +\frac{1}{N}\sum_{l,m}{l\left[\frac{1}{2}\left(\delta_{lm}+\delta_{l,m+1}\right)\gamma\left\langle\hat{d}_{m}^{\dagger}\hat{d}_{m}\right\rangle\right]}$
$\ -\frac{1}{N}\sum_{l,m}{il\left\langle\hat{n}_{l}\right\rangle\left\langle\hat{H}_{\text{eff}}^{\dagger}-\hat{H}_{\text{eff}}\right\rangle}$
$\ +\frac{1}{N}\sum_{l,m}{l\gamma\left\langle\hat{d}_{m}^{\dagger}\hat{c}_{l}\right\rangle\left\langle\hat{c}_{l}^{\dagger}\hat{d}_{m}\right\rangle.}$
Appendix B Feedback protocol with boundary effect for the tilted potential
In this section, we focus on a different feedback protocol, considering the boundary effects with PBCs. This is because a tilted chain exhibits peculiar phenomena under PBCs, behaving similarly to OBCs and showing anomalous bulk dynamics. We incorporate the edge, considering the edge feedback protocol, and the quantum jump operator is written as,
$\hat{L}_{l}^{\prime}=\begin{cases}e^{i\theta\hat{n}_{l+1}}\hat{d}_{l}^{\dagger}\hat{d}_{l}&\text{ if }l=1,2,...,L-1,\\ e^{i\theta\hat{n}_{1}}\hat{d}_{L}^{\dagger}\hat{d}_{L}&\text{ if }l=L,\end{cases}$ (14)
where $\hat{d}_{L}^{\dagger}=(\hat{c}_{L}^{\dagger}-i\hat{c}_{1}^{\dagger})/\sqrt{2}$. The edge feedback protocol (hereafter referred to as “edge feedback”) differs from the previous feedback protocol, which will be abbreviated as “previous feedback” in this section. The quantum jump with the edge feedback accounts for the additional “left moving” from the initial site to the final one.
From FIG. 6 (a) and (b), $\overline{\left\langle\hat{n}_{l}\right\rangle}$ for previous and edge feedback are the same without the tilted potential. The bipartite entanglement entropy $\overline{S}_{L/2}$ are the same for different feedback likewise, from green solid lines in FIG. 7. Comparing FIG. 6 (c) and (d), despite some similarities, the steady state in (d) is not completely occupied in the left part, and the right part shows a sparse density distribution. It is evident from the orange solid lines in FIG. 7 that the $\overline{S}_{L/2}$ for the edge feedback is consistently higher than the other. The reason for the phenomenon is that edge feedback affects the pseudo-edge effect, which is caused by the tilted potential. A minority of particles can cross the pseudo edge from the first to the last site with the aid of the quantum jump with the edge feedback. If $\Delta=0$, the absence of the pseudo edge occurs, so observables remain the same for different feedback.